answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

When tossing 2 coins the probability of getting exactly 1 tail is one over 3?

No, the possible results of two coin flips are;

HH

HT

TH

TT

The probability of getting exactly one tail is 1/2 or 50%.

How many ways can you choose 2 books from 5 on a shelf?

You can wind up with 10 different pairs of books in your hand, which

you can choose from a shelf of 5 books in 20 different ways.

What are the names of the queens in a deck of cards?

Queen of Hearts, Queen of Diamonds, Queen of Spades, Queen of Clubs.

What is the or rule in probability?

Given two events, A and B, the probability of A or B is the probability of occurrence of only A, or only B or both.

In mathematical terms: Prob(A or B) = Prob(A) + Prob(B) - Prob(A and B).

What is the probability of three digit numbers without repeated digits with numbers 0123456?

It is possible to create a 3-digit number, without repeated digits so the probability is 1.

How many permutation of umbrella?

There are 20160 permutations of the word.

What is the probability of first rolling a 5 and then rolling a 3 using a fair six-sided die?

Probability of ' 5 ' on the first roll = 1/6

Probability of ' 3 ' on the second roll = 1/6

Probability of ' 5 ' followed by ' 3 ' = 1/6 x 1/6 = 1/36 = 27/9 percent

What percent of data falls between 1 Standard deviation below and 2 stand deviations above the mean?

The answer will depend on what the distribution is. Non-statisticians often assum that the variable that they are interested in follows the Standard Normal distribution. This assumption must be justified.

If that is the case then the answer is 81.9%

What is the probability of rolling and odd number on one toss of a number cube?

On a typical 6-sided number cube, there are three odd numbers (1, 3, 5) and three even numbers (2, 4, 6). Therefore, there are three outcomes that are desired out of a total of six possible outcomes. 3/6 = 1/2, so your probability is 1 in 2 or 50%.

What is the probability that exactly two times head are being tossed?

The answer depends on how many times the coin is tossed. The probability is zero if the coin is tossed only once!

Making some assumptions and rewording your question as "If I toss a fair coin twice, what is the probability it comes up heads both times" then the probability of it being heads on any given toss is 0.5, and the probability of it being heads on both tosses is 0.5 x 0.5 = 0.25.

If you toss it three times and want to know what the probability of it being heads exactly twice is, then the calculation is more complicated, but it comes out to 0.375.

How many ways can you rearrange small?

The letter is the word "small" can be rearranged in 60 different ways.

What is the probability that a couple will have 4 boys and 1 girl?

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a boy is approx 0.52, the probability 4 boys and 1 girl out of 5 children is 0.1724 approx.

What is the probability of drawing a 6 or diamond from a deck of 52 cards?

The answer depends on how many cards are drawn and whether they are drawn at random.

The probability that a single card, drawn at random from a deck of cards in 16/52 = 4/13.

What is the difference between a simple random sample and a stratified random sample?

Sometimes a population consists of a number of subsets (strata) such that members within any particular strata are alike while difference between strata are more than simply random variations. In such a case, the population can be split up into strata. Then a stratified random sample consists of simple random samples, with the same sampling proportion, taken within each stratum.

Trending Questions
How are probabilities of independent events alike? What dice game has dice with multicolored sides two yellow two red and two green roll ten dice and similar to farkle? What is the probability that from a normal 52 deck of card you randomly draw a three and then without replacing the3 you draw the queen of hearts? The random variable Z is the standard normal score. Find the number a if P Za equals 0.9830? What is the probability of drawing a card from each of 4 decks and getting the nine of clubs from all decks? What is the probabiliy of selecting odd or prime numbers from 1 to 50? A fair coin is tossed 3 times what is the probability it will land on heads all 3 tosses? What is the sample space if you roll 2 dice? Suppose you roll two number cubes once Use the counting principle to find the probability of rolling a 4 on both cubes? If the variance of a distribution increases will the confidence in a given interval decrease? What is the probability of getting an even number if the die is rolled once? Does a king and jack beat a king and 4 high cards? Is there a being one 5? What is the probability of drawing a king and a jack in two successive draws from a deck of cards with replacement? What are theoretical events? True or false the calculation of probabilities is based on planned events? How can probabilities be used in genetics? How ways cam you arrange a 4 letter word Adam? What is the probability of getting an odd number when rolling a single 6-sided die? Is a standard deviation dependent on the mean?